Given: A -3.85% grade meets a -2.90% grade at station 41+25 (VPI). The EVC elevation 227.49 ft. The length is 550’. Calculate the full (+00) stations on the curve.
Answer #1

C OE 100 J.-9, = 2.Gut 3.85 +0.952 Ele (PVC) = Ele(Eve) - (9+) = 227.49-(-2.9–3,85 x 550 ce Eve Xo = 246.05 Sta(PVC)= No = s